is -54 an answer to the cubed root of 4x+11 = 5?
Answer by mathrebel(25)

solve:
(4*x+51)^(1/3)=5
first, eliminate the cube root by raising the equation to the third power:
(4*x+51)=5^3
4*x+51=125
subtract the 25 from both sides:
4x=74
then divide by 4 both sides
x=18.5
-54 is wrong????